# Short vs. Long-term Coordination of Drones: When Distributed Optimization Meets Deep Reinforcement Learning

### v5

@@ III-B Problem Formulation
Mission efficiency. It denotes the ratio of sensing values in all cells at all timeslots collected by the drones during their mission over the total required values in all cells at all timeslots during the period $t$. It is formulated as:
@@
Energy cost. It is the energy consumed by drones to perform spatio-temporal sensing. A power consumption model [33] is used to calculate the power consumption with input the specification of drones (weight, propeller and power efficiency). This model estimates the cost of navigation and sensing plans, and emulates the outdoor environments [16]. The energy cost of each drone $u$ is formulated as:
$E(a^{u},t)=C^{\mathsf{f}}(u)\cdot t^{\mathsf{f}}(a^{u})+C^{\mathsf{h}}(u)\cdot t^{\mathsf{h}}(a^{u}),$ (5)
@@
where $C^{\mathsf{f}}(u)$ and $C^{\mathsf{h}}(u)$ are the flying and hovering power consumption of drone $u$ respectively; $t^{\mathsf{f}}$ and $t^{\mathsf{h}}$ are the flying and hovering time respectively, which are determined by the sensing plan $P^{u}(a^{u},t)$.
@@
Overall performance. The optimization problem has three objectives: 1) maximize mission efficiency; 2) maximize sensing accuracy; and 3) minimize energy cost. Simultaneously achieving all of these objectives is a daunting task, as they often oppose each other. Maximizing both efficiency and accuracy is a trade-off since increasing efficiency of mission results in sensing imbalance and even blind areas by under-sensing [25]. In addition, in order to achieve both of these goals, drones have to continuously move to collect data from different cells, which results in higher energy costs due to long-distance flights. However, some trips of drones may not improve sensing quality in the long run. For instance, at the end of a period, drones opt to return to the nearest charging station where there are fewer sensing tasks in the next period. This choice reduces the energy cost of the current period but comes at the significant expense of sacrificing the sensing quality in the next period. Therefore, the overall sensing performance is optimized by integrating these opposing objectives and selecting the plans of each drone. The objective function is formulated as follows:
$\mathop{\arg\max}\limits_{a^{u},u\in\mathcal{U}}\;\sum^{U}_{u=1}\sum^{T}_{t=1}[\mathsf{Eff}(a^{u},t)\cdot\mathsf{Acc}(a^{u},t)-E(a^{u},t)].$ (6)
DRL Modeling. Once drones take actions of flying directions, they execute plans and return to charging stations, changing their current locations and battery levels while observing the navigation and sensing of other drones. Therefore, the problem scenario can be explained into a Markov decision process [26]. We model the problem using state, action, and reward concepts: (i) State: The state $\textbf{s}_{t}$ at period $t$ consists of four components $(\mathcal{S}_{1},\mathcal{S}_{2},\mathcal{S}_{3},\mathcal{S}_{4})$, where $\mathcal{S}_{1}=\{c^{u}(t)|_{u\in\mathcal{U}}\}$ is the current locations of drones. Since drones charge at charging stations before taking actions, the location can serve as the index of the charging station. $\mathcal{S}_{2}=\{e^{u}(t)|_{u\in\mathcal{U}}\}$ is the current battery levels of drones, which are calculated based on the battery capacity and energy cost. $\mathcal{S}_{3}=\{P^{u}(t)|_{u\in\mathcal{U}}\}$ is the plan of $u$. $\mathcal{S}_{4}=\{P^{-u}(t)|_{u\in\mathcal{U}}\}$ is the aggregated plan of other drones excluding $u$, which are shared via the designed communication model. (ii) Action: The action $\textbf{a}_{t}=\{a^{1}_{t},...,a^{U}_{t}\}$ at period $t$ consists of period-by-period flying directions (N, E, S, W, NE, SE, SW, NW or O) determined by drones. (iii) Reward: Based on the objective function of Eq.(6), the expected immediate local reward of one drone at period $t$ is defined as follows: $\textbf{r}^{u}_{t}=\mathsf{Eff}(a^{u},t)\cdot\mathsf{Acc}(a^{u},t)-E(a^{u},t)$. Throughout the training process or episodes, the overall reward fluctuates based on the actions taken by the drones. This helps drones in prioritizing areas rich in sensor data by maximizing their reward, i.e., the highest cumulative overall performance.
Figure 1: Overview of the structured tree communication model. During the bottom-up process, Drone 3 aggregates the plans of its children, i.e., drone 1 and drone 2, and sends them to its parent agent with its own plan. During the top-down process, each parent agent sends the total aggregated plans to its children such that all agents obtain the observed plans.
III-C Communication Model
In DRL model, each drone evaluates an immediate reward by considering collective information from other drones. To enhance information exchange efficiency, we design a structured tree communication model. Each drone, controlled by a local agent, connects with other agents into a tree communication structure, as shown in Fig. 1, within which it interacts with its children and parent in a bottom-up and top-down. For example, drones aggregate and share their plans $P^{u}$ in a bottom-up fashion, and then obtain the observed plans $P^{-u}$ in the top-down process [8]. Therefore, this decentralized coordination among agents is highly-efficient with low communication cost [8]. Through iterative bottom-up and top-down exchanges, agents can flexibly change their plans based on the aggregated navigation and sensing from other drones.
Furthermore, agents can be self-organized within the tree to make the system more resilient to attacks or failures [12, 11]. Since agents share the aggregated states instead of the agents’ one, the system is scalable and designed to preserve privacy [^2] Additional privacy protection mechanisms such as differential privacy and homomorphic encryption apply here [34].. For instance, assuming a swarm of drones that belong to various entities, such as aviation agencies, private citizens, and companies, they collaborate to create shared navigation and sensing strategies which rely on the collective sum of plans rather than individual private sensing information.
@@ III-C Communication Model
III-D System Overview
For high-quality sensing performance, the proposed approach DO-RL relies on both decentralized drone coordination for short-term navigation and sensing optimization, as well as long-term scheduling of flying directions to adapt to changing environments. Fig.2 illustrates the designed system framework of DO-RL, consisting of four main components:
Plan generation. This component generates the navigation and sensing options for drones. Given the sensing map, each drone autonomously generates a finite number of discrete plans to initialize the overall process. This provides flexibility for the drones at the next stage to choose in a coordinated way. The generated plans are grouped into different flying directions. This initialization stage is executed only once for each drone.
@@ III-D System Overview
DRL-based scheduling. This component is the core of the framework, which leverages the DRL algorithm to enable drones to take their period-by-period actions of flying directions. Then, the generated subset of plans with the corresponding navigation action are chosen as input for the plan selection. After plan selection, it undergoes two steps: one is periodic state update, which updates the state of drones (i.e., the current location and energy cost) and the system (i.e., the target and required sensing values) for the DRL-based scheduling in the next period; the other is multi-agent reinforcement learning, which is built based on centralized training and decentralized execution.
IV Detailed System Design
@@
IV-A Plan Generation
Each drone generates nine groups of plans, each associated with an action for a flying direction. These groups consist of multiple individual plans of the same size. A group of plans generated by drone $u$ during a period is defined as $G^{a^{u}}(t)$, where each plan in the group has the same action, $P^{u}(a^{u},t)\in G^{a^{u}}(t)$. Then, a plan can be defined as $P^{u}(a^{u},t)=\{l,K(a^{u}),J(a^{u}),E(a^{u},t)\}$, where $l$ is the index of the plan, $l\leq L$, ($L$ denotes the total number of plans in a group); $K(a^{u})$ denotes the indexes of cells within the searching range along flying direction $a^{u}$; $J(a^{u})$ denotes the indexes of visited cells within $K(a^{u})$; and $E(a^{u},t)$ denotes the energy consumed by $u$ traveling over $J(a^{u})$, i.e., the cost of the plan.
A drone generates a plan by determining its path and calculating the energy required for traversing that path, as shown in Fig. 3. As an example of a plan, the drone $u$ firstly finds a searching range to cells $K(a^{u})$: it determines a range with a certain width (equal to half of distance to nearest charging station) along its flying direction, and then searches the cells within this range. Next, it selects the visited cells $J(a^{u})$ randomly from this range. Then, $u$ finds the shortest possible path among $J(a^{u})$ via the Dijkstra’s algorithm for the Traveling Salesmen Problem. The path includes the information of which cell and which timeslot the drone uses to collect data, and thus both the traveling time $t^{\mathsf{f}}(a^{u})$ and hovering time $t^{\mathsf{h}}(a^{u})$ are calculated. The total energy consumption $E(a^{u},t)$ when $u$ travels over the path is then calculated via Eq.(5).
IV-B Plan Selection
Given the plans in selected flying directions within a period, drones can adapt their assignments by selecting appropriate plans in response to changes in task targets, enhancing their flexibility and adaptability in dynamic environments. In this process, the agents of drones improve their plan selections based on the structured tree communication model. Specifically, each agent obtains the aggregated choices, i.e., the aggregated plan $P^{-u}(a^{u},t)$, from other agents, and chooses one of the plans $P^{u}(a^{u},t,l)$ such that all choices together $P^{u}(a^{u},t)+P^{-u}(a^{u},t)$ add up to match a target $\mathcal{R}(t)$. The target is used to steer each drone to sense over a unique cell during a timeslot, thereby avoiding the simultaneous sensing of multiple drones within the same cell, i.e., preventing over-sensing and under-sensing.
@@ IV-B Plan Selection
IV-C Periodic State Update
After choosing a plan, each drone changes its current state at the next period $t+1$, including its location $c^{u}(t+1)$, battery level $e^{u}(t+1)$, selected plan $P^{u}(t+1):=P^{u}(a^{u},t)$, and observed plans $P^{-u}(t+1):=P^{-u}(a^{u},t)$. However, due to dynamic changing environment, drones have no knowledge about the required sensing value in the next period. They need to predict and estimate the required sensing value to calculate reward function via Eq.(3) and (4). Depending on the predicted distribution of sensor data, the target is required to be updated.
@@ IV-C Periodic State Update
where $V^{\prime}_{n,s}$ denotes the data values collected by drones once they execute their selected plans (the total collected values are aggregated and shared to each agent via the top-down interactions in the distributed optimization process [8]); $\omega_{n,s}(t^{\prime})$ is a prediction coefficient such that $10<\omega_{n,s}(t^{\prime})<$ 1, $t^{\prime}\leq t$. To achieve accurate predictions, DO-RL leverages the Ordinary Least Squares regression method (OLS) to train these coefficients initially, and use the past experienced observations as the target distribution for training. After that, the current collected data values $V^{\prime}_{n,s}$ and predicted one $\hat{V}_{n,s}$ are shared to all agents via the tree communication structure.
@@
IV-D Multi-agent Reinforcement Learning
Our approach provides a generalized model for multi-agent reinforcement learning by adopting the framework of centralized training and decentralized execution, as shown in Fig.4. During execution, each agent obtains a local aggregated observation comprising both private and public information. Apart from their own private data such as locations, battery levels, and selected plans, this observation contains aggregated plans shared via the designed communication model (shown in Fig. 1). The training part is built based on an actor-critic model. The actors allow drones to work independently and choose actions based on their observations and policies, allowing for parallel data collection and processing. The critic, however, evaluates the actions of drones and guide them toward collectively optimal decisions. In addition, DO-RL employs Proximal Policy Optimization (PPO) to prevent detrimental updates and improving the stability of the learning process [35].
@@ V-A Experimental Settings
Scenario model. In order to evaluate the DO-RL, we model a real-world transportation scenario, where a swarm of drones perform the sensing tasks of traffic monitoring. The number of vehicles serves as the required sensing value in the model. Drones rely on accurate vehicle observation to detect early signs of traffic congestion. This enables traffic operators to apply mitigation measures, reducing the carbon footprint in one of the world’s highest-emission sectors.. The scenario involves a real-world map of Munich city, imported to the simulations of urban mobility (SUMO) [^3] https://www.eclipse.org/sumo/ to accurately generate realistic flows of vehicles. Fig. 5 illustrates a selected map of $1600\times 1600$ meters in the city with the simulation time of $50$ hours, which is $100$ periods. It has a high density of vehicles, approximately $2,000$ vehicles passing by per hour. The area split into a finite number of cells, each is defined as a rectangular square with the size of $200\times 200$ meters that can be captured by the cameras of drones. Fig. 6 shows the distribution of both charging stations and traffic vehicles in the map. The charging stations are uniformly distributed in the map. We employ cross-validation: $80\%$ simulation time of the datasets for training and $20\%$ for testing. There are two scenarios for the experimental evaluation in this paper: (i) Basic scenario. It has $64$ cells, $4$ charging stations and $8$ time periods (set as $30min$ for each period). It has high density of vehicles, around $2,000$ vehicles passing by per time period. $16$ drones sense the area (camera recording) in parallel. (ii) Complex scenario. It varies the parameter settings such as the density of drones and vehicles, as well as the number of time periods, cells and charging stations.
@@
Drones. We assume that all drones are of the same type, specifically the DJI Phantom 4 Pro model [^4] https://www.dji.com/uk/phantom-4-pro/info. Each drone has a body weight of $1.07$ kg and four propellers, each with a diameter of $0.35$ m. The ground speed of the drones is set at $6.94$ m/s, and the drag force is determined to be $4.1134$[33]. As a consequence, the power consumption remains consistent across all drones. Additionally, drones are equipped with a $6000$ mAh LiPo 2S battery ($0.31$ kg) [36]. The maximum flying time for drones is approximately $30$ min, which is set as one flying period. A flying period has $S=30$ timeslots, each of one minute. Moreover, to ensure that the camera covers the entire area of a cell (see Fig. 5), the minimum hovering height of drones is determined at which the field of view of the camera and the cell overlap. The hovering height $h$ is calculated based on the distance between any two cells (approximately $200$ meters), the camera resolution, focal length derived from camera calibration, and ground sampling distance [37]. Therefore, each drone equipped with a 4K camera senses from a minimum height of $82.4$ meters.
@@
Planning and distributed optimization. We leverage the decentralized multi-agent collective learning method of EPOS [^5] EPOS is open-source and available at: https://github.com/epournaras/EPOS., the Economic Planning and Optimized Selections [8]. We generate $L=64$ plans for each agent. All generated plans are made openly available to encourage further research on coordinated spatio-temporal sensing of drones [16, 25, 13]. Each agent in EPOS is mapped to a drone. During the coordinated plan selection via EPOS, agents self-organize into a balanced binary tree as a way of structuring efficient learning interactions [38]. The shared goal of the agents is to minimize the RMSE between the total sensing values collected and the target via Eq.(7). In the one execution of EPOS, the agents perform $40$ bottom-up and top-down learning iterations during which RMSE converges to a minimum optimized value. To minimize the RMSE and the cost of plans, we make an empirical choice of parameters, such as the number of visited cells (mobility range) $J^{u}=2$ and the behavior of agents $\beta$ within a range $[0.1,0.8]$. The results are shown in Fig. 7.
@@
Neural network and learning algorithm. We sample $H=64$ groups of transitions as minibatches in a replay buffer, and set the discount factor as $\gamma=0.95$ and the clip interval hyperparameter to $0.2$ for policy updating. Then we try different multi-layer perceptions (MLP), and use $W=64$ neurons in the three hidden layers of the MLP in both critic and actor networks. The activation function used for the networks is tanh. We train the models for $\mathcal{E}=5000$ episodes with multiple epochs (i.e., time period).
@@
Complexity. Given the number of nodes per layer $W$, number of episodes $\mathcal{E}$, size of state space $\mathcal{X}$, and number of actions $A$, the computational complexity of DNN is approximately $O(C_{dnn}):=O((\mathcal{X}+A)W+3W^{2})$ [18]. The comparison of both computational and communication cost is shown in Table III, where $I$ denotes the number of iterations in EPOS.
TABLE III: Comparison of computational and communication costs.

### v6

@@ III-B Problem Formulation
Mission efficiency. It denotes the ratio of sensing values in all cells at all timeslots collected by the drones during their mission over the total required values in all cells at all timeslots during the period $t$. The purpose is to collect sensing data as much as possible. It is formulated as:
@@
However, maximizing this metric leads to sensing imbalance and even blind areas. Some cells may be covered for a long time while some other cells may be never covered, i.e., the over-sensing and under-sensing [24]. Thus, it is necessary to consider matching indicators such as the sensing accuracy [24] to access such imbalances.
@@
Apart from improving the effciency and accuracy of drone sensing, the energy consumed by drones needs to be saved.
@@
Energy cost. It is the battery usage of drones to perform spatio-temporal sensing. A power consumption model [36] is used to calculate the power consumption with input the specification of drones (weight, propeller and power efficiency). This model estimates the cost of navigation and sensing plans, and emulates the outdoor environments [15]. The energy cost of each drone $u$ is formulated as:
$E(a^{u},t)=\frac{C^{\mathsf{f}}(u)\cdot t^{\mathsf{f}}(a^{u})+C^{\mathsf{h}}(u)\cdot t^{\mathsf{h}}(a^{u})}{B(u)},$ (5)
@@
where $B(u)$ is the battery capacity of drone $u$; $C^{\mathsf{f}}(u)$ and $C^{\mathsf{h}}(u)$ denote the flying and hovering power consumption of $u$ respectively; $t^{\mathsf{f}}$ and $t^{\mathsf{h}}$ are the flying and hovering time respectively, which are determined by the sensing plan $P^{u}(a^{u},t)$.
@@
As described before, the optimization problem has three objectives: 1) maximize mission efficiency; 2) maximize sensing accuracy; and 3) minimize energy cost. Simultaneously achieving all of these objectives is a daunting task, as they often oppose each other. Maximizing both efficiency and accuracy is a trade-off. In addition, in order to achieve both of these goals, drones have to continuously move to collect data from different cells, which results in higher energy costs due to long-distance flights. However, some trips of drones may not improve sensing quality in the long run. For instance, at the end of a period, drones opt to return to the nearest charging station where there are fewer sensing tasks in the next period. This choice reduces the energy cost of the current period but comes at the significant expense of sacrificing the sensing quality in the next period. Therefore, the overall sensing performance is optimized by integrating these opposing objectives and selecting the plans of each drone.
Overall performance. The objective function is defined as the weighted sum of mission efficiency, sensing accuracy and energy cost:
$\begin{split}\mathop{\arg\max}\limits_{a^{u},u\in\mathcal{U}}\sum^{U}_{u=1}\sum^{T}_{t=1}[\;&\alpha_{1}\mathsf{Eff}(a^{u},t)+\alpha_{2}\mathsf{Acc}(a^{u},t)\\ &-\alpha_{3}E(a^{u},t)].\end{split}$ (6)
IV The Proposed Solution
@@
IV-A System Overview
In our problem, traditional DRL approaches suffer from the large action space and dimensional curse caused by the increasing number of time periods, leading to decline in sensing quality. To ensure high-quality sensing performance, the proposed DO-RL integrates decentralized drone coordination for short-term navigation and sensing optimization with long-term scheduling of flying directions. Fig.1 illustrates the designed system framework of DO-RL, consisting of three main components:
Plan generation. This component generates the navigation and sensing options for drones. Given the sensing map, each drone autonomously generates a finite number of discrete plans to initialize the overall process. This provides flexibility for the drones at the next stage to choose in a coordinated way. The generated plans are grouped into different flying directions.
@@ IV-A System Overview
DRL-based scheduling. This component is the core of the framework, which leverages the DRL algorithm to enable drones to take the actions of overall flying directions between departure and destination charging stations in each time period. Thus, these actions are executed period-by-period, with each made only after the drone completes its sensing missions in the current period. The component includes two elements: a periodic state update, which updates the state of drones for DRL-based scheduling in the next period, and a multi-agent reinforcement learning module, which is built based on centralized training and decentralized execution.
The DO-RL process works as follows: Each drone initially generates a set of plans. It then selects a flying direction and chooses a subset of plans aligned with that direction. Each drone autonomously picks a plan from this subset, shares it with the swarm, and observes others’ choices. Based on its action and observations, the drone calculates a reward function and updates its state (including location, battery level, and sensing requirements), and stores these results in a multi-agent reinforcement learning buffer to refine its flight strategies. This cycle of DRL-based scheduling and plan selection repeats at each time period until the mission is complete. However, DO-RL requires frequent information sharing, which makes communication inefficient and vulnerable to potential failure of individual drones.
Figure 2: Overview of the structured tree communication model. During the bottom-up process, Drone 3 aggregates the plans of its children, i.e., drone 1 and drone 2, and sends them to its parent agent with its own plan. During the top-down process, each parent agent sends the total aggregated plans to its children such that all agents obtain the observed plan.
IV-B Communication Model
To overcome these barriers, DO-RL leverages a structured tree communication model. Each drone, controlled by a local agent, locates and connects with other agents into a tree communication structure, as shown in Fig. 2. Once agents have established their proximity-based relationship, e.g., Euclidean distance, they are positioned starting from the leaves up to the root, each interacting with its children and parent in a bottom-up and top-down [11, 10]. For example, a drone $u$ aggregates and shares its plan $P^{u}$ in a bottom-up fashion, and then obtains the observed plan $P^{-u}$ in the top-down process [7]. These plans serve as the observation of an agent in DRL modeling. As this coordination among agents using hierarchical organization reduces the need for every node to communicate with every other node, the system is highly-efficient in collaboration and information exchange [7].
Through iterative bottom-up and top-down exchanges, agents can flexibly change their plans and adapt to dynamic sensing environments. This ensures that even if some nodes fail, the overall sensing performance remains effective. Moreover, the system can dynamically reconfigure, repositioning drones within the tree in response to communication failures, ensuring that a single drone failure has minimal impact on information flow. The resilience of this model has been previously demonstrated in [11] with large-scale real-world datasets, and thus this paper focuses on sensing performance and scalability of the proposed solution.
V Detailed System Design
V-A DRL Modeling
The core design of DO-RL lies in applying DRL on our problem. Once drones take actions of flying directions, they execute plans and return to charging stations, changing their current locations and battery levels while observing the navigation and sensing of other drones. Therefore, the problem scenario can be explained into a Markov decision process [6]. We model the problem using state, action, and reward concepts: (i) State: The state $\textbf{s}_{t}$ at period $t$ consists of four components $(\mathcal{S}_{1},\mathcal{S}_{2},\mathcal{S}_{3},\mathcal{S}_{4})$, where $\mathcal{S}_{1}=\{c^{u}(t)|_{u\in\mathcal{U}}\}$ is the current locations of drones. Since drones charge at charging stations before taking actions, the location can serve as the index of the charging station. $\mathcal{S}_{2}=\{e^{u}(t)|_{u\in\mathcal{U}}\}$ is the current battery levels of drones, which are calculated based on the energy cost. $\mathcal{S}_{3}=\{P^{u}(t)|_{u\in\mathcal{U}}\}$ is the plan of $u$. $\mathcal{S}_{4}=\{P^{-u}(t)|_{u\in\mathcal{U}}\}$ is the aggregated plan of other drones excluding $u$, which are shared via the structured communication model. (ii) Action: The action $\textbf{a}_{t}=\{a^{1}_{t},...,a^{U}_{t}\}$ at period $t$ consists of period-by-period flying directions (N, E, S, W, NE, SE, SW, NW or O) determined by drones. (iii) Reward: Based on the objective function of Eq.(6), the expected immediate local reward of one drone at period $t$ is defined as follows: $\textbf{r}^{u}_{t}=\alpha_{1}\mathsf{Eff}(a^{u},t)+\alpha_{2}\cdot\mathsf{Acc}(a^{u},t)-\alpha_{3}E(a^{u},t)$. Throughout the training process or episodes, the overall reward fluctuates based on the actions taken by the drones. This helps drones in prioritizing areas rich in sensor data by maximizing their reward, i.e., the highest cumulative overall performance.
@@ V-A DRL Modeling
V-B Plan Generation
The actions enable each drone to generate nine groups of plans, each associated with an action for a flying direction. These groups consist of multiple individual plans of the same size. A group of plans generated by drone $u$ during a period is defined as $G^{a^{u}}(t)$, where each plan in the group has the same action, $P^{u}(a^{u},t)\in G^{a^{u}}(t)$. Then, a plan can be defined as $P^{u}(a^{u},t)=\{l,K(a^{u}),J(a^{u}),E(a^{u},t)\}$, where $l$ is the index of the plan, $l\leq L$, ($L$ denotes the total number of plans in a group); $K(a^{u})$ denotes the indexes of cells within the searching range along flying direction $a^{u}$; $J(a^{u})$ denotes the indexes of visited cells within $K(a^{u})$; and $E(a^{u},t)$ denotes the energy consumed by $u$ traveling over $J(a^{u})$, i.e., the cost of the plan.
The details of generating a single plan are defined by the chosen path and the corresponding energy required to traverse that path, as shown in Fig. 3. As an example of a plan, the drone $u$ firstly finds a searching range to cells $K(a^{u})$: it determines a range with a certain width (equal to half of distance to nearest charging station) along its flying direction, and then searches the cells within this range. Next, it selects the visited cells $J(a^{u})$ randomly from this range. Then, $u$ finds the shortest possible path among $J(a^{u})$ via the Dijkstra’s algorithm for the Traveling Salesmen Problem. The path includes the information of which cell and which timeslot the drone uses to collect data, and thus both the traveling time $t^{\mathsf{f}}(a^{u})$ and sensing time $t^{\mathsf{h}}(a^{u})$ are calculated. The total energy consumption $E(a^{u},t)$ when $u$ travels over the path is then calculated via Eq.(5).
V-C Plan Selection
Given the plans in selected flying directions within a period, drones can adapt their assignments by selecting appropriate plans in response to changes in task targets, enhancing their adaptability in dynamic environments. In this process, the agents of drones improve their plan selections via tree communication. Specifically, each agent obtains the aggregated choices, i.e., the aggregated plan $P^{-u}(a^{u},t)$, from other agents, and chooses one of the plans $P^{u}(a^{u},t,l)$ such that all choices together $P^{u}(a^{u},t)+P^{-u}(a^{u},t)$ add up to match a target $\mathcal{R}(t)$. The target is used to steer each drone to sense over a unique cell during a timeslot, thereby avoiding the simultaneous sensing of multiple drones within the same cell, i.e., preventing over-sensing and under-sensing.
@@ V-C Plan Selection
V-D Periodic State Update
After choosing a plan, each drone changes its current state at the next period $t+1$, including its location $c^{u}(t+1)$, battery level $e^{u}(t+1)$, selected plan $P^{u}(t+1):=P^{u}(a^{u},t)$, and observed plan $P^{-u}(t+1):=P^{-u}(a^{u},t)$. However, due to dynamic changing environment, drones have no knowledge about the required sensing value in the next period. They need to predict and estimate the required sensing value to calculate reward function via Eq.(3) and (4). Depending on the predicted distribution of sensor data, the target is required to be updated.
@@ V-D Periodic State Update
where $V^{\prime}_{n,s}$ denotes the data values collected by drones once they execute their selected plans; $\omega_{n,s}(t^{\prime})$ is a prediction coefficient such that $10<\omega_{n,s}(t^{\prime})<$ 1, $t^{\prime}\leq t$. To achieve accurate predictions, DO-RL leverages the Ordinary Least Squares regression method (OLS) to train these coefficients initially, and use the past experienced observations as the target distribution for training.
@@
At the end of state update in DO-RL, the current collected data values $V^{\prime}_{n,s}$, the predicted values $\hat{V}_{n,s}$ and the target $\mathcal{R}$ are shared with each agent via the top-down interactions within the tree communication structure. This information will be stored and later sampled for multi-agent reinforcement learning.
V-E Multi-agent Reinforcement Learning
Our approach provides a generalized model for multi-agent reinforcement learning by adopting the framework of centralized training and decentralized execution, as shown in Fig.4. During execution, each agent obtains a local aggregated observation comprising both private and public information. Apart from their own private data such as locations, battery levels, and selected plans, this observation contains aggregated plans shared via the tree communication model (shown in Fig. 2). The training part is built based on an actor-critic model. The actors allow drones to work independently and choose actions based on their observations and policies, allowing for parallel data collection and processing. The critic, however, evaluates the actions of drones and guide them toward collectively optimal decisions. In addition, DO-RL employs Proximal Policy Optimization (PPO) to prevent detrimental updates and improving the stability of the learning process [37].
@@ VI-A Experimental Settings
Scenario model. In order to evaluate the DO-RL, we model a real-world transportation scenario, where a swarm of drones perform the sensing tasks of traffic monitoring. The number of vehicles serves as the required sensing value in the model. The scenario involves a real-world map of Munich city, imported to the simulations of urban mobility (SUMO) [^2] https://www.eclipse.org/sumo/ to accurately generate realistic flows of vehicles. Fig. 5 illustrates a selected map of $1600\times 1600$ meters in the city with the simulation time of $50$ hours, which is $100$ periods. It has a high density of vehicles, approximately $2,000$ vehicles passing by per hour. The area split into a finite number of cells, each is defined as a rectangular square with the size of $200\times 200$ meters that can be captured by the cameras of drones. Fig. 6 shows the distribution of both charging stations and traffic vehicles in the map. The charging stations are uniformly distributed in the map. We employ cross-validation: $80\%$ simulation time of the datasets for training and $20\%$ for testing. There are two scenarios for the experimental evaluation in this paper: (i) Basic scenario. It has $64$ cells, $4$ charging stations and $8$ time periods (set as $30min$ for each period). It has high density of vehicles, around $2,000$ vehicles passing by per time period. $16$ drones sense the area (camera recording) in parallel. (ii) Complex scenario. It varies the parameter settings such as the density of drones and vehicles, as well as the number of time periods, cells and charging stations. Noted that the simulated traffic flow can be displayed in a big screen in a testbed experimentation using hardware to emulate the complex outdoor sensing environments, which has been proved by our previous work in [15, 34].
@@
Drones. We use DJI Phantom 4 Pro model [^3] https://www.dji.com/uk/phantom-4-pro/info. Each drone has a body weight of $1.07$ kg and four propellers, each with a diameter of $0.35$ m. The ground speed of the drones is set at $6.94$ m/s, and the drag force is determined to be $4.1134$[36]. As a consequence, the power consumption remains consistent across all drones. Additionally, drones are equipped with a $6000$ mAh LiPo 2S battery ($0.31$ kg). The maximum flying time for drones is approximately $30$ min, which is set as one flying period. A flying period has $S=30$ timeslots, each of one minute. Moreover, to ensure that the camera covers the entire area of a cell (see Fig. 5), the minimum hovering height of drones is determined at which the field of view of the camera and the cell overlap. The hovering height $h$ is calculated based on the distance between any two cells (approximately $200$ meters), the camera resolution, focal length derived from camera calibration, and ground sampling distance [38]. Therefore, each drone equipped with a 4K camera senses from a minimum height of $82.4$ meters.
@@
Planning and distributed optimization. We leverage the decentralized multi-agent collective learning method of EPOS [^4] EPOS is open-source and available at: https://github.com/epournaras/EPOS., the Economic Planning and Optimized Selections [7]. We generate $L=64$ plans for each agent. All generated plans are made openly available to encourage further research on coordinated spatio-temporal sensing of drones [15, 24, 12]. Each agent in EPOS is mapped to a drone. During the coordinated plan selection via EPOS, agents self-organize into a balanced binary tree as a way of structuring efficient learning interactions [39]. The shared goal of the agents is to minimize the RMSE between the total sensing values collected and the target via Eq.(7). In the one execution of EPOS, the agents perform $40$ bottom-up and top-down learning iterations during which RMSE converges to a minimum optimized value.
@@
Neural network and learning algorithm. In our problem, the weight parameters $\alpha_{1}$, $\alpha_{2}$ and $\alpha_{3}$ are all set to a value of $1$. A total of $H=64$ groups of transitions are sampled as minibatches in a replay buffer, with a discount factor of $\gamma=0.95$ and a clip interval hyperparameter of $0.2$ for policy updating. We try the recurrent neural network (RNN), and use $W=64$ neurons in the three hidden layers of the RNN in both critic and actor networks. The activation function used for the networks is tanh. The models are trained over $\mathcal{E}=5000$ episodes, each consisting of multiple epochs (equal to the number of time periods).
@@
Complexity. Given the number of nodes per layer $W$, number of episodes $\mathcal{E}$, size of state space $\mathcal{X}$, and number of actions $A$, the computational complexity of DNN is approximately $O(C_{dnn}):=O((\mathcal{X}+A)W+3W^{2})$ [25]. The comparison of both computational and communication cost is shown in Table II, where $I$ denotes the number of iterations in EPOS. The results illustrate that DO-RL significantly lowers the computation cost by reducing the number of actions required during training on a temporal scale, compared to the state-of-the-art MAPPO. DO-RL also has low communication cost due to the tree communication structure.
TABLE II: Comparison of computational and communication costs.